OpenStudy (anonymous):

A Web music store offers two versions of a popular song. The size of the standard version is 2.8 megabytes (MB). The size of the high-quality version is 4.1 MB. Yesterday, the high-quality version was downloaded four times as often as the standard version. The total size downloaded for the two versions was 4032 MB. How many downloads of the standard version were there?

OpenStudy (anonymous):

@DDCamp can you help me when your done please?

OpenStudy (ddcamp):

Let S be the downloads of the standard version and H be the downloads of the high quality version. What do we know about H and S? 2.8S+4.1H = 4032 (total magabytes downloaded) H = 4S (we're told the high quality was downloaded four times as often)

OpenStudy (theviper):

Solve it by using Substitution method :- H=4S (It is given) 2.8S+4.1H=4032 As we know 'H=4S' Therefore, 2.8S+4.1(4S)=4032

OpenStudy (anonymous):

i got it thanks

OpenStudy (theviper):

welcome